OpenStudy (anonymous):

Identify the values of a, b, and c in the following quadratic equation. 4x^2 + x + 9 = 0

OpenStudy (anonymous):

a = , b = , c =

Parth (parthkohli):

\( \color{Black}{\Rightarrow ax^2 + bx + c =0 }\)

Parth (parthkohli):

I gave you the standard equation. Compare and contrast.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

options for each 4 9 1

OpenStudy (anonymous):

4x^2 + (1)x + 9 = 0 ax^2 + bx + c = 0 compare and contrast further, hehe

OpenStudy (anonymous):

how

OpenStudy (anonymous):

well a,b, and c are the coefficients. A coefficient, meaning, a number placed in front of a variable (like x). All you are asked to do, is identify the numbers in front of the "x"'s. And in the right order. From left to right, a,b,c, as the equation goes.

OpenStudy (anonymous):

So when you look at the two equations right next to each other, it should be fairly obvious which number is a, which number is b, and which is c. As they are almost directly next to each other when I posted them 2 posts ago

OpenStudy (anonymous):

ok